### v1

# The Tournament Theorem of Rédei revisited

Abstract
In 1934 L. Rédei published his famous theorem that the number of Hamiltonian paths in a tournament is odd. In fact it is a corollary of a stronger theorem in his paper. Stronger theorems were also obtained in the early 1970s by G.A. Dirac in his lectures at Aarhus University and by C. Berge in his monographs on graphs and hypergraphs. We exhibit the stronger theorems of Rédei, Dirac and Berge and explain connections between them. The stronger theorem of Dirac has two corollaries, one equivalent to Rédei’s stronger theorem and the other related to Berge’s stronger theorem.
AMS Subject Classification: 05C20
Keywords: Digraphs, Tournaments, Hamilton Path
1 Introduction
When Beineke, Toft and Wilson worked on Milestones in Graph Theory [1] it was noticed that the 1933 paper by L. Rédei [4] contains a stronger theorem than Rédei’s Theorem that a tournament always has an odd number of Hamiltonian paths:
Theorem 1.1.
(Rédei’s Stronger Theorem) Let $T$ be a tournament with at least $2$ vertices. Add to $T$ a new non-empty set $W$ of vertices and some new non-oriented edges between $W$ and $T$ and between vertices of $W$. Then the number of Hamiltonian paths in the new graph, beginning and ending in $T$, is even.
Note that 0 is an even number! A hamiltonian path in the new graph contains a non-empty set of non-oriented edges and a set (perhaps empty) of oriented edges, oriented in the correct forward direction of the Hamiltonian path. In case the Hamiltonian path consists of only non-oriented edges then the reverse path is counted as a second Hamiltonian path (this statement may be avoided by having oriented double-edges, one in each direction, replacing non-oriented edges – in fact this is how Rédei formulated his stronger theorem. We shall overcome the difficulty by using permutation, rather than Hamiltonian paths).
If $W$ consists of just one vertex joined to exactly two vertices $x$ and $y$ of $T$ then Rédei’s Stronger Theorem implies that reversing the direction of the edge between $x$ and $y$ in $T$ results in a new tournament $T^{\prime}$, where the parity of the number of Hamiltonian paths in $T$ and $T^{\prime}$ are the same. Since the transitive tournament has exactly one Hamiltonian path it follows that $T$ has an odd number of Hamiltonian paths.
In [2] and in its later editions and translations C. Berge deduced Rédei’s Theorem from a stronger result (where again a Hamilton path with only non-oriented edges is counted twice):
Theorem 1.2.
(Berge’s Stronger Theorem) Let $G$ be a mixed graph with at least $2$ vertices and let $\mkern 1.5mu\overline{\mkern-1.5muG\mkern-1.5mu}\mkern 1.5mu$ be its complement (an undirected edge and a non-edge are complements; an edge oriented from $x$ to $y$ and the oriented edge from $y$ to $x$ are complements). Then the numbers of Hamiltonian paths in $G$ and $\mkern 1.5mu\overline{\mkern-1.5muG\mkern-1.5mu}\mkern 1.5mu$ have the same parity - both even or both odd (both cases occur).
Asking Carsten Thomassen whether Rédei’s Stronger Theorem is generally known, he informed us that G.A. Dirac in lectures at Aarhus University in the early 1970s also proved stronger versions of Rédei’s Theorem. Consulting Dirac’s handwritten lecture notes [3] we have generalized his approach and proved what we shall call Dirac’s Stronger Theorem. It has a special case equivalent to Rédei’s Stronger Theorem. One attractive corollary of Dirac’s Stronger Theorem is the following result (where again a Hamilton path with only non-oriented edges is counted twice):
Theorem 1.3.
(Dirac’s Corollary 1) Let $G$ be a complete mixed graph on at least $2$ vertices (each pair of vertices is joined by either a non-oriented edge or by an oriented edge). Then the number of Hamiltonian paths with all oriented edges in the correct direction containing at least one non-oriented is even.
In the following we shall state and prove Dirac’s Stronger Theorem and obtain relations to the stronger theorems of Rédei and Berge.
2 Dirac’s Stronger Theorem
aabbcc
Figure 1: A mixed graph $G$ with $n=3$ vertices.
In a mixed graph $G$ each pair of vertices constitutes either a non-edge, a non-oriented edge or an oriented edge in one of the two possible directions. We shall denote the set of non-edges in $G$ by $E_{1}$, the set of non-oriented edges in $G$ by $E_{2}$ and the set of oriented edges in $G$ by $E_{3}$. Moreover, by $\mkern 1.5mu\overline{\mkern-1.5muE_{3}\mkern-1.5mu}\mkern 1.5mu$ we shall denote the set $E_{3}$ with all orientations reversed. A permutation $\mathbf{x}$ of $G$ is an ordering of the vertices of the mixed graph $G$ on $n$ vertices ($n\geq 2$):
$\mathbf{x}=(x_{1},x_{2},...,x_{i},x_{i+1}...,x_{n}).$
For the neighboring pair $e=x_{i}x_{i+1}$ of the permutation $\mathbf{x}$ we have the following possibilities. Either $e=\{x_{i},x_{i+1}\}$ belongs to $E_{1}\cup E_{2}$ (i.e., $e$ is a non-edge or a non-oriented edge of $G$), or $e=(x_{i},x_{i+1})$ belongs to $E_{3}\cup\mkern 1.5mu\overline{\mkern-1.5muE_{3}\mkern-1.5mu}\mkern 1.5mu$ (i.e., $e\in E_{3}$ is an oriented edge of $G$ and has the correct direction or $e\in\mkern 1.5mu\overline{\mkern-1.5muE_{3}\mkern-1.5mu}\mkern 1.5mu$ is an oriented edge of $\mkern 1.5mu\overline{\mkern-1.5muG\mkern-1.5mu}\mkern 1.5mu$ and has the wrong direction. We denote by $E(\mathbf{x})$ the set of neighboring pairs of $\mathbf{x}$. An element $e\in E_{1}\cup E_{2}\cup E_{3}\cup\mkern 1.5mu\overline{\mkern-1.5muE_{3}\mkern-1.5mu}\mkern 1.5mu$ is contained in $\mathbf{x}$ if $e\in E(\mathbf{x})$. Furthermore, we denote by $\mathbb{P}(G)$ the set of permutations of $G$. Note that if $G$ has $n\geq 2$ vertices, then we have $|E_{1}|+|E_{2}|+|E_{3}|=\binom{n}{2}$, $|\mathbb{P}(G)|=n!$ and $|E(\mathbf{x})|=n-1$ for every $\mathbf{x}\in\mathbb{P}(G)$. A Hamilton path of $G$ is a permutation $\mathbf{x}\in\mathbb{P}(G)$ such that $E(\mathbf{x})\subseteq E_{2}\cup E_{3}$.
Let $G$ be the mixed graph with $n=3$ vertices depicted in Figure 1. Then $E_{1}=\{\{b,c\}\}$, $E_{2}=\{\{a,b\}\}$, $E_{3}=\{(a,c)\}$, and $\mkern 1.5mu\overline{\mkern-1.5muE_{3}\mkern-1.5mu}\mkern 1.5mu=\{(c,a)\}$. For the permutation $\mathbf{x}=(c,a,b)\in\mathbb{P}(G)$, we have $E(\mathbf{x})=\{(c,a),\{a,b\}\}=\mkern 1.5mu\overline{\mkern-1.5muE_{3}\mkern-1.5mu}\mkern 1.5mu\cup E_{2}$. If $\mathbb{P}^{\prime}=\{\mathbf{x}\in\mathbb{P}(G)\;|\;\{a,b\}\in E(\mathbf{x})\}$, then we have $\mathbb{P}^{\prime}=\{\mathbf{x}^{1}=(a,b,c),\mathbf{x}^{2}=(b,a,c),\mathbf{x}^{3}=(c,a,b),\mathbf{x}^{4}=(c,b,a)\}$. The only Hamilton path of $G$ is the permutation $\mathbf{x}=(b,a,c)$.
With the above notation we can now state:
Theorem 2.1.
(Dirac’s Stronger Theorem) Let $G$ be a mixed graph with $n\geq 2$ vertices. Let $A$ be a subset of $E_{1}\cup E_{2}$, Let $N_{A}$ denote the number of permutations of $G$ containing all elements of $A$ and no element from $\mkern 1.5mu\overline{\mkern-1.5muE_{3}\mkern-1.5mu}\mkern 1.5mu$. Let $N_{=A}$ denote the number of permutations of $G$ containing precisely $A$ as its elements from $E_{1}\cup E_{2}$ and containing no element from $\mkern 1.5mu\overline{\mkern-1.5muE_{3}\mkern-1.5mu}\mkern 1.5mu$. Then $N_{A}$ and $N_{=A}$ have the same parity. In particular $N_{A}-N_{=A}$ is even.
Dirac’s Corollary 1 follows from Dirac’s Stronger Theorem with $A$ and $E_{1}$ both empty. To prove the theorem we shall use the following lemma (with the notation from Dirac’s Stronger Theorem):
Lemma 2.2.
Let $G$ be a mixed graph with $n\geq 2$ vertices. Let $A$ be a subset of $E_{1}\cup E_{2}$ and let $D$ be a subset of $\mkern 1.5mu\overline{\mkern-1.5muE_{3}\mkern-1.5mu}\mkern 1.5mu$. Let $N$ be the number of permutations of $G$ containing all elements of $A\cup D$. Then $N$ is even, except when |A|+|D|=n1|A|+|D|=n-1, $|D|\geq 1$, and $A\cup D=E(\mathbf{x})$ for a permutation $\mathbf{x}\in\mathbb{P}(G)$, in which case $N=1$.
Proof.
Clearly, $N=0$ except perhaps when $A\cup D$ forms the elements of disjoint sub-permutations of $G$, say $p$ sub-permutations without any element from $D$ and $q$ sub-permutations with at least one element from $D$. In particular $|A|+|D|\leq n-1$. If we also consider each single vertex, not incident with any element of $A\cup D$, as a sub-permutation, then let the total number of sub-permutations be $r$ (in fact r=n|A||D|r=n-|A|-|D|). Then
$N=r!\cdot 2^{p}.$
It follows that $N$ is even, except when $r=1$, $p=0$, and hence |A|+|D|=n1|A|+|D|=n-1, $|D|\geq 1$, and $A\cup D=E(\mathbf{x})$ for a permutation $\mathbf{x}\in\mathbb{P}(G)$. In this exceptional case $N=1$. This proves the lemma. ∎
Proof of Theorem 2.1.
A permutation of $G$ contains n1n-1 elements. Hence, if $|A|\geq n$, then $N_{A}=N_{=A}=0$ and if |A|=n1|A|=n-1, then $N_{A}=N_{=A}.$ So assume that $|A|\leq n-2$. We may also assume that $A$ forms the elements of $p$ disjoint sub-permutations, for otherwise we have $N_{A}=N_{=A}=0$. By Lemma 2.2, for every set $D\subseteq\mkern 1.5mu\overline{\mkern-1.5muE_{3}\mkern-1.5mu}\mkern 1.5mu$, the number $M(D)$ of permutations of $G$ containing $A\cup D$ is even, except when |D|=n1|A||D|=n-1-|A| and $A\cup D=E(\mathbf{x})$ for a permutation $\mathbf{x}\in\mathbb{P}(G)$, in which case $M(D)=1$. Since $N_{A}$ is the number of permutations of $G$ containing $A$ and no element of $\mkern 1.5mu\overline{\mkern-1.5muE_{3}\mkern-1.5mu}\mkern 1.5mu$, the inclusion-exclusion principle leads to
$N_{A}=\sum_{D\subseteq\mkern 1.5mu\overline{\mkern-1.5muE_{3}\mkern-1.5mu}\mkern 1.5mu,|D|\leq n-1-|A|}(-1)^{|D|}M(D)$
Consequently, modulo 2, $N_{A}$ is congruent to the number of sets $D$ of n1|A|n-1-|A| elements from $\mkern 1.5mu\overline{\mkern-1.5muE_{3}\mkern-1.5mu}\mkern 1.5mu$ that together with $A$ forms the elements of a permutation of $G$; this number is equal to $N_{=A}$ (by reversing the orders of the permutations). Hence, we have $N_{A}\equiv N_{=A}\,\mathrm{(mod}\;2\mathrm{)}$. ∎
The two cases of Theorem 2.1, where $A=\varnothing$ or $A=(E_{1}\cup E_{2})\setminus\{e\}$ for an element $e\in E_{1}\cup E_{2}$, imply the following two corollaries:
Theorem 2.3.
(Dirac’s Corollary 2) Let $G$ be a mixed graph with $n\geq 2$ vertices. Let $N_{\varnothing}$ denote the number of permutations $\mathbf{x}\in\mathbb{P}(G)$ satisfying $E(\mathbf{x})\subseteq E_{1}\cup E_{2}\cup E_{3}$. Let $N_{=\varnothing}$ denote the number of permutations containing no elements from $E_{1}\cup E_{2}\cup\mkern 1.5mu\overline{\mkern-1.5muE_{3}\mkern-1.5mu}\mkern 1.5mu$. Then $N_{\varnothing}$ and $N_{=\varnothing}$ have the same parity. Hence $N_{\varnothing}-N_{=\varnothing}$ is even, i.e., the number of permutations containing at least one element from $E_{1}\cup E_{2}$, but no element from $\mkern 1.5mu\overline{\mkern-1.5muE_{3}\mkern-1.5mu}\mkern 1.5mu$ is even.
Theorem 2.4.
(Dirac’s Corollary 3) Let $G$ be a mixed graph with $n\geq 2$ vertices and with $E_{1}\cup E_{2}\neq\varnothing$, let $e\in E_{1}\cup E_{2}$, and let $A=(E_{1}\cup E_{2})\setminus\{e\}$. Then $N_{A}$ and $N_{=A}$ have the same parity. Hence $N_{A}-N_{=A}$ is even, i.e. the number of permutations containing all elements from $E_{1}\cup E_{2}$, but no element from $\mkern 1.5mu\overline{\mkern-1.5muE_{3}\mkern-1.5mu}\mkern 1.5mu$ is even.
3 Rédei’s Stronger Theorem
Consider the situation in Rédei’s Stronger Theorem (Theorem 1.1) with a tournament $T$, an extra non-empty set of vertices $W$ and extra non-oriented edges between vertices in $W$ and between $W$ and $T$. Let the new mixed graph be denoted $G$. Let $\cal P$} be a set of paths in $G$, each path joining two vertices in $T$ and with all its interior vertices from $W$, such that each vertex in $W$ belongs to exactly one path and such that the paths of $\cal P$} together form a set of perhaps fewer disjoint paths (with vertices of $T$ possibly as interior). If no such set of paths exists then the number of Hamiltonian paths in $G$, starting and ending in $T$, is zero and hence even. For each path from $\cal P$} replace the oriented edge in $T$ joining its endvertices by an unoriented edge. Then Dirac’s Corollary 3 (Theorem 2.4) implies that the number of Hamiltonian paths in $G$, starting and ending in $T$, is even. Summation over all possible pathsystems $\cal P$} produces thus an even number. This shows that Dirac’s Corollary 3 implies Rédei’s Stronger Theorem.
Consider conversely the situation in Dirac’s Corollary 3. For each element from $E_{1}\cup E_{2}$ add an extra new vertex joined to the two vertices of the element by two non-oriented edges. Let the set of new vertices be $W$. Then Rédei’s Stronger Theorem implies Dirac’s Corollary 3.
The above shows that Rédei’s Stronger Theorem is equivalent with Dirac’s Corollary 3.
4 Berge’s Stronger Theorem
Let $G$ be a mixed graph as before. Moreover, let $\mathbb{P}_{0}$ denote the permutations without elements from $\mkern 1.5mu\overline{\mkern-1.5muE_{3}\mkern-1.5mu}\mkern 1.5mu$, let $\mathbb{P}_{i}$ the permutations without elements from $E_{i}\cup\mkern 1.5mu\overline{\mkern-1.5muE_{3}\mkern-1.5mu}\mkern 1.5mu$ for $i=1$ and $2$, and let $\mathbb{P}_{3}$ the permutations without elements from $E_{1}\cup E_{2}\cup\mkern 1.5mu\overline{\mkern-1.5muE_{3}\mkern-1.5mu}\mkern 1.5mu$ (in other words $\mathbb{P}_{3}=\mathbb{P}_{1}\cap\mathbb{P}_{2})$; see Figure 2.
$\mathbb{P}_{0}:~\mkern 1.5mu\overline{\mkern-1.5muE_{3}\mkern-1.5mu}\mkern 1.5mu\text{ - free}$ $\mathbb{P}_{1}:(E_{1}\cup\mkern 1.5mu\overline{\mkern-1.5muE_{3}\mkern-1.5mu}\mkern 1.5mu)\text{ - free}$ $\mathbb{P}_{2}:(E_{2}\cup\mkern 1.5mu\overline{\mkern-1.5muE_{3}\mkern-1.5mu}\mkern 1.5mu)\text{ - free}$ $\mathbb{P}_{3}=\mathbb{P}_{1}\cap\mathbb{P}_{2}$
Figure 2: The sets $\mathbb{P}_{i}\subseteq\mathbb{P}(G)$ with $i\in\{0,1,2,3\}$.
With this notation Dirac’s Corollary 2 (Theorem 2.3) is the statement that $|\mathbb{P}_{0}|-|\mathbb{P}_{3}|$ is even, and hence that $|\mathbb{P}_{0}|+|\mathbb{P}_{3}|$ is even. Berge’s Stronger Theorem (Theorem 1.2) is that $|\mathbb{P}_{1}|+|\mathbb{P}_{2}|$ is even (note that the number of permutations without elements from $E_{2}\cup E_{3}$ equals $|\mathbb{P}_{2}|$ by reversing the orders of the permutations). We do not see an easy argument getting from one of these two results to the other. But note that if $G$ is complete (that is $E_{1}=\varnothing$) then $|\mathbb{P}_{0}|=|\mathbb{P}_{1}|$ and $|\mathbb{P}_{2}|=|\mathbb{P}_{3}|$, and Dirac’s result and Berge’s result are the same. Similarly, when $E_{2}=\varnothing$.
The above results of Dirac and Berge together imply the following theorem:
Theorem 4.1.
(Berge-Dirac Theorem) Let $G$ be a mixed graph with $n\geq 2$ vertices. Let $\mathbb{P}_{0}$ denote the set of permutations of $G$ without elements from $\mkern 1.5mu\overline{\mkern-1.5muE_{3}\mkern-1.5mu}\mkern 1.5mu$, let $\mathbb{P}_{i}$ denote the set of permutations of $G$ without elements from $E_{i}\cup\mkern 1.5mu\overline{\mkern-1.5muE_{3}\mkern-1.5mu}\mkern 1.5mu$ $(i\in\{1,2\})$, and let $\mathbb{P}_{3}=\mathbb{P}_{1}\cap\mathbb{P}_{2}$. Then:
$|\mathbb{P}_{0}|+|\mathbb{P}_{1}|+|\mathbb{P}_{2}|+|\mathbb{P}_{3}|\mbox{ is even.}$
We do not know any simpler proof of the Berge-Dirac Theorem than the above combination. Assuming the Berge-Dirac Theorem it follows that Dirac’s Corollary 2 and Berge’s Stronger Theorem are equivalent. Also note that in the Berge-Dirac Theorem we may replace any + by a -, for example
$|\mathbb{P}_{0}|-|\mathbb{P}_{1}|-|\mathbb{P}_{2}|+|\mathbb{P}_{3}|$ is even
which means that the number of permutations containing at least one element from $E_{1}$, at least one element from $E_{2}$, and no element from $\mkern 1.5mu\overline{\mkern-1.5muE_{3}\mkern-1.5mu}\mkern 1.5mu$ is even.